OpenStudy (anonymous):

Help me please! Somebody, ANYBODY ! -2x+8y=18 -2x+y=11 how do i do this?

OpenStudy (dumbcow):

Elimination method. eliminate the x variable by subtracting equations -2x - (-2x) = 0 8y - y = 7y 18 - 11 = 7 -> 7y = 7 y = 1 substitute back in to find x_value

OpenStudy (anonymous):

Substitute y from the 2nd equation into the first eq. y = 2x+11 -2x+8(2x+11) = 18 -2x + 16x + 88 = 18 14x = -70 x = -5 y = -10+11 y = 1

OpenStudy (anonymous):

thank you! but, now how do i get x ?

OpenStudy (dumbcow):

Or take the value of y and substitute it back into original equation

OpenStudy (dumbcow):

giorgia already worked it out for you...